OpenStudy (anonymous):

write the slope - intercept form of the equation of the line passing through the point (-6,-2) and parallel to the line y=4x-1

OpenStudy (texaschic101):

y = 4x - 1 the slope here is 4. A parallel line will have the same slope. Now we will use y = mx + b slope(m) = 4 (-6,-2)...x = -6 and y = -2 now we sub, and since we already know the slope(m), we need to find the y intercept (b). -2 = 4(-6) + b -2 = -24 + b -2 + 24 = b 22 = b so your equation in slope intercept form is : y = 4x + 22